# Knowledge-guided machine learning model with soil moisture for corn yield prediction under drought conditions

### v1

@@ Abstract
Remote sensing (RS) techniques, by enabling non-contact acquisition of extensive ground observations, have become a valuable tool for corn yield prediction. Traditional process-based (PB) models are limited by fixed input features and struggle to incorporate large volumes of RS data. In contrast, machine learning (ML) models are often criticized for being “black boxes” with limited interpretability. To address these limitations, we used Knowledge-Guided Machine Learning (KGML), which combined the strengths of both approaches and fully used RS data. However, previous KGML methods overlooked the crucial role of soil moisture in plant growth. To bridge this gap, we proposed the Knowledge-Guided Machine Learning with Soil Moisture (KGML-SM) framework, using soil moisture as an intermediate variable to emphasize its key role in plant development. Additionally, based on the prior knowledge that the model may overestimate under drought conditions, we designed a drought-aware loss function that penalizes predicted yield in drought-affected areas. Our experiments showed that the KGML-SM model outperformed other ML models. Finally, we explored the relationships between drought, soil moisture, and corn yield prediction, assessing the importance of various features and analyzing how soil moisture impacts corn yield predictions across different regions and time periods.
@@
Corn yield prediction; Knowledge-guided Machine learning; Soil moisture; drought
@@
[inst2]organization=Department of of Soil Science,addressline=University of Wisconsin-Madison, city=Madison, postcode=53706, state=WI, country=USA
@@ 1 Introduction
Corn plays a vital place in the U.S. as a primary crop, supporting food security as well as the animal feed and biofuel industries (Graham et al., 2007; Thompson, 1969). Accurate yield prediction is crucial for effective resource management and economic stability (Kucharik and Ramankutty, 2005). However, achieving precise prediction across large areas and diverse environmental conditions is challenging, making it necessary to pursue advanced research in corn yield prediction techniques (Weiss et al., 2020). Traditional process-based (PB) models (Weiss et al., 2007) are built on a deep understanding of the physical, chemical, and biological processes within crop growth systems (Weiss et al., 2004). These models use equations derived from scientific principles to simulate the behavior of crops, weather, and soil interactions (Weiss et al., 2002), making them highly interpretable and reliable in scenarios where the underlying processes are well understood (Puntel et al., 2016; Shahhosseini et al., 2021). Several promising modeling frameworks are designed for simulating and predicting agricultural production systems, such as The Agricultural Production Systems sIMulator (APSIM) (McCown et al., 1996), Decision Support System for Agrotechnology Transfer (DSSAT) (Jones et al., 2003), WOrld FOod STudies (WOFOST) (Van Diepen et al., 1989), and The Agricultural Policy / Environmental eXtender (APEX) (Williams and Izaurralde, 2010). Many recent studies (Zhen et al., 2023, 2022) have successfully used these models to predict crop yields. However, PB models have several limitations. Plant growth is highly complex, and these models rely on a limited set of fixed input features, which are insufficient to accurately capture the intricacies of this process. Moreover, PB models require extensive manual parameter tuning, limiting their applicability for large-scale applications across vast regions.
@@
In recent years, remote sensing (RS) has provided several benefits for corn yield prediction (Lobell et al., 2015). It allows for the collection of large-scale and real-time data across vast agricultural areas, which, in conjunction with machine learning (ML) methods, provides an accurate and efficient way for large-scale yield prediction. Deep learning (Goodfellow et al., 2016), a widely used ML method, has the advantage of leveraging vast amounts of RS data by automatically learning features from large datasets. This removes the need for manual feature engineering, enabling the model to capture complex patterns and relationships between environmental factors and crop yield (He et al., 2016). Many studies have applied deep learning to crop yield prediction and shown its advantages. You et al. (2017) introduced a scalable approach to crop yield prediction using modern representation learning rather than traditional hand-crafted features, achieving impressive results. Wang et al. (2018) demonstrated soybean yield predictions in Argentina and obtained reliable results in Brazil using transfer learning, even with limited data. Ma et al. (2021a) developed a county-level corn yield prediction model using a Bayesian Neural Network across 12 states in the Corn Belt, showing the potential for large-scale corn yield prediction tasks. However, despite the popularity of deep learning models in corn yield prediction tasks, they have certain limitations. First, deep learning models require a large amount of data to learn the relationships between inputs and outputs. Second, deep learning models often function as a black box, lacking reliable interpretability of the prediction results. Therefore, finding ways to expand our data sources is a critical factor in achieving accurate predictions. Additionally, there are complex biological processes among soil, weather, and corn, which are challenging for deep learning models to fully capture. Given these challenges, combining PB models with deep learning models has become an important research topic.
@@
Knowledge-guided machine learning (KGML) (Karpatne et al., 2022) aims to integrate scientific knowledge in ML frameworks to achieve better generalizability, scientific consistency, and explainability of results. KGML involves three main approaches: Knowledge-Guided Learning, Knowledge-Guided Architecture, and Knowledge-Guided Pretraining. Knowledge-Guided Learning incorporates scientific knowledge, such as physical laws, into the learning algorithm by modifying the loss function. It guides the optimization process toward scientifically consistent and generalizable solutions (Daw et al., 2022; Bao et al., 2021). Knowledge-Guided Architecture embeds scientific knowledge directly into the structure of the ML model. By designing knowledge-guided neural network architectures, researchers can enable models to automatically capture physical properties or conservation laws, which improves both interpretability and robustness (Dugdale et al., 2017; Luo et al., 2023). Knowledge-Guided Pretraining uses simulated data or self-supervised learning tasks to pretrain ML models. By initializing model parameters using data that encodes scientific principles, this method reduces the need for large amounts of observational data and improves performance in data-scarce scenarios (Licheng et al., 2022; Chen et al., 2023). Building on these approaches, KGML has been increasingly applied in various domains, including agriculture, hydrology, and climate science, enabling improved data-driven decision-making and predictive modeling.
@@
Recently, many studies have used KGML to combine PB models and ML models for crop yield prediction (Kimball et al., 2023; Burroughs et al., 2023). Since PB models already provide insights into the crop growth process and generate high-quality simulated data, Knowledge-Guided Pretraining is the most popular approach (He et al., 2023; Yang et al., 2023). These studies aim to use ML to identify patterns between crop features and yield from large datasets, while also deriving interpretability from the data simulated by PB models. Specifically, KGML models leverage PB models to simulate intermediate variables representing key plant growth processes, enhancing their ability to capture domain-specific knowledge. Then, ML models are applied to simulate specific parts of the process. By incorporating these intermediate variables, the model can better represent the underlying biological and environmental mechanisms. For example, He et al. (2023) integrated physical knowledge from PB models and enabled the model to adjust for temporal data shifts over the years. Yang et al. (2023) used both historical and in-season RS data to better predict key agricultural variables such as grain yield and carbon cycling processes. Yang et al. (2024) presented a knowledge-guided computer vision framework designed to monitor and simulate the growth of individual strawberry fruits, which can predict growth at the individual fruit level by integrating image-based data and dynamic simulations.
Although these KGML models have made significant progress in crop yield prediction, all existing KGML-based studies on crop yield prediction have overlooked the critical role of soil, particularly how soil moisture contributes to plant growth and affects yield. Soil moisture is influenced by temperature, precipitation, and other environmental factors and directly affects crop growth. Unlike RS data, soil moisture has a cumulative impact on yield over time. As a key drought indicator, low soil moisture causes water stress, which limits corn growth and reduces yields (Unganai and Kogan, 1998). Nonetheless, both PB and ML models have certain limitations in modeling soil moisture. In PB models, weather data serve as input, while soil moisture is treated as output (McCown et al., 1996; Williams and Izaurralde, 2010), failing to capture its direct impact on crop yield. In ML models, a key challenge is the limited soil moisture data. For example, as the most commonly used soil moisture dataset, the SPL4SMGP.007 SMAP dataset (Entekhabi et al., 2010) only covers data from 2015 onward, limiting ML models’ ability to learn from extensive historical data and impacting the model performance. Therefore, we integrated PB and ML models through KGML to better model soil moisture for corn yield prediction. Additionally, we examined how soil moisture variations over time and across space affected corn yield in drought-prone areas, where drought often reduced yields and hindered prediction accuracy (Cooper et al., 2014). Researchers have long been interested in studying the relationship between soil moisture, drought, and corn yield prediction (Ines et al., 2013; Bushong et al., 2016; Mladenova et al., 2017). However, these studies either have a narrow scope (Vergopolan et al., 2021), resulting in limited data that cannot be effectively generalized to broader applications, or they focus solely on end-of-season yield prediction (Pignotti et al., 2023).
In this paper, we proposed the KGML-SM framework, which integrated PB and ML models for corn yield prediction while incorporating the role of soil moisture in plant growth. The framework consisted of two key components: a Weather-to-Soil (W2S) encoder that modeled the relationship between weather inputs and soil moisture outputs, and an attention module (Vaswani, 2017) that integrated features from multiple sources and dynamically weighted their influence on corn yield across different growth stages and drought conditions. Additionally, we designed a drought-aware yield prediction loss function, informed by the observed relationship between drought and yield reduction, to further enhance model performance under drought conditions. To train this model, we adopted a two-step process. First, we used the APSIM model to generate simulated soil moisture and corn yield data, which were then combined with the input weather data for pretraining. Then, we finetuned the model with soil moisture data from the SMAP product, corn yield data from official statistics, as well as weather data and vegetation indices (VIs) from Google Earth Engine (GEE) (Gorelick et al., 2017), which enabled it to adapt to real-world variability and enhance predictive accuracy. Our research aimed to improve the accuracy and interpretability of corn yield predictions, which ultimately contributed to more resilient agricultural systems in the face of climate variability.
@@ 2 Data acquisition
In this study, we developed two datasets for corn yield prediction. The first, a field-level dataset, was generated using APSIM (McCown et al., 1996) and used for pretraining. The second, a county-level dataset, was derived from GEE (Gorelick et al., 2017) and USDA NASS and employed for finetuning. The workflow was divided into three main steps: simulation, pretraining, and finetuning. A summary of all public datasets used in this study is provided in Table 1. In this section, we first introduce the study area (Sec. 2.1); then, we provide details of the APSIM field-level dataset (Sec. 2.2); finally, we describe the construction of the GEE county-level dataset (Sec. 2.3).
@@
MODIS c Vegetation indices (Justice et al., 1998; Schaaf and Wang, 2015)
@@
SMAP c Soil moisture data (Entekhabi et al., 2010)
@@ 2.1 Study area
Our research focused on corn yield prediction across the U.S. Corn Belt, selecting twelve states as our study area: North Dakota, South Dakota, Minnesota, Wisconsin, Iowa, Illinois, Indiana, Ohio, Missouri, Kansas, Nebraska, and Michigan. These states are crucial agricultural regions in the U.S., known for their significant contributions to corn production. We generated a five-year average yield map (Fig. 1) for these twelve states and considered them well suited for corn yield prediction research.
@@ 2.2 APSIM field-level dataset
In this section, we explain how APSIM was used to generate a field-level simulated dataset for model pretraining. For the simulation of the APSIM field-level dataset, we used data from the Iowa Environmental Mesonet (IEM) (Herzmann et al., 2004), a platform developed by Iowa State University that provides U.S. agricultural and environmental data. We downloaded all available station data from the IEM, which included data from thousands of stations across 12 states from 1980 to 2023 (Table 2).
@@
State IA IL IN KS MI MN MO ND NE OH SD WI
Station number 113 120 89 150 140 140 130 79 128 100 114 147
2.2.1 Input and output of APSIM
The APSIM model simulates crop yield using four weather data inputs: maximum temperature (Tmax), minimum temperature (Tmin), precipitation (PPT), and radiation (Radn). It also requires certain management parameters, which we describe in detail later (Sec. 3.3.1). Maximum and minimum temperatures influence the rate of plant development and growth stages, impacting processes like photosynthesis and respiration. Precipitation data are essential for modeling soil moisture levels, which directly affect water availability for crops. Radiation is a critical factor in photosynthesis, as it provides the energy needed for plant growth. These weather data are available from the IEM website (Herzmann et al., 2004), which collects and distributes environmental data from various observing networks. The APSIM model takes weather data as input to simulate root zone soil moisture (SM_rootzone), surface soil moisture (SM_surface), and corn yield. SM_rootzone is crucial for corn growth as it directly affects water availability for uptake, influencing plant development and yield. SM_surface plays a key role in seed germination and early growth stages.
2.2.2 Field-level dataset for pretraining
Besides APSIM’s input and output variables, we also extracted two useful features from IEM: prediction year and location (longitude and latitude). Then, we calculated the five-year historical average yield based on the average local simulated corn yield over the past five years. We believed these three features might contribute to improving the accuracy of our corn yield prediction model. After APSIM generated the simulation data, we combined its input weather data, output soil moisture, and corn yield, along with other features to construct the APSIM field-level dataset (Table 3). This dataset was used for pretraining our KGML model, aiming to help the model learn the internal processes of corn growth.
Table 3: Summary of variables comprising the APSIM field-level dataset, including inputs, outputs, and other features

### v2

# Knowledge-guided machine learning for county-level corn yield prediction under drought

@@ Abstract
Remote sensing (RS) technique, enabling the non-contact acquisition of extensive ground observations, is a valuable tool for crop yield predictions. Traditional process-based models struggle to incorporate large volumes of RS data, and most users lack understanding of crop growth mechanisms. In contrast, machine learning (ML) models are often criticized as “black boxes” due to their limited interpretability. To address these limitations, we utilized Knowledge-Guided Machine Learning (KGML), a framework that leverages the strengths of both process-based and ML models. Existing works have either overlooked the role of soil moisture in corn growth or did not embed this effect into their models. To bridge this gap, we developed the Knowledge-Guided Machine Learning with Soil Moisture (KGML-SM) framework, treating soil moisture as an intermediate variable in corn growth to emphasize its key role in plant development. Additionally, based on the prior knowledge that the model may overestimate under drought conditions, we designed a drought-aware loss function that penalized predicted yield in drought-affected areas. Our experiments showed that the KGML-SM model outperformed other traditional ML models. We explored the relationships between drought, soil moisture, and corn yield prediction by assessing the importance of different features within the model, and analyzing how soil moisture impacts predictions across different regions and time periods. Finally we provided interpretability for prediction errors to guide future model optimization.
@@
Corn yield prediction; Knowledge-guided Machine learning; Soil moisture; drought-aware learning
@@
[inst2]organization=Department of of Soil and Environmental Sciences,addressline=University of Wisconsin-Madison, city=Madison, postcode=53706, state=WI, country=USA
@@ 1 Introduction
Corn, as a primary crop, plays a vital role in the U.S. agriculture, supporting food security as well as the animal feed and biofuel industries (Graham et al., 2007; Thompson, 1969). Accurate yield prediction is crucial for effective resource management and economic stability (Kucharik and Ramankutty, 2005). Achieving accurate prediction across large areas and diverse environmental conditions is challenging (Weiss et al., 2020). Traditional process-based models (Weiss et al., 2007) are built on a deep understanding of the physical, chemical, and biological processes within crop growth systems (Weiss et al., 2004). These models use equations derived from scientific principles to simulate the behavior of crops, weather, and soil interactions (Weiss et al., 2002), making them highly interpretable and reliable in scenarios where the underlying processes are well understood (Puntel et al., 2016; Shahhosseini et al., 2021). Several promising modeling frameworks are designed for simulating and predicting agricultural production systems, such as the Agricultural Production Systems sIMulator (APSIM) (McCown et al., 1996), the Decision Support System for Agrotechnology Transfer (DSSAT) (Jones et al., 2003), the WOrld FOod STudies (WOFOST) (Van Diepen et al., 1989), and the Agricultural Policy / Environmental eXtender (APEX) (Williams and Izaurralde, 2010). Many studies (Zhen et al., 2023, 2022) have successfully used these models to predict crop yields. However, process-based models have several limitations. Plant growth is highly complex, and these models rely on a limited set of fixed input features, which are insufficient to accurately capture the intricacies of this process. Moreover, process-based models require extensive manual parameter tuning, limiting their applicability for applications across vast regions.
@@
In recent years, remote sensing (RS) has provided several benefits for corn yield prediction (Lobell et al., 2015). It enables the collection of large-scale real-time data across vast agricultural areas. This data, typically combined with machine learning (ML), facilitates accurate and efficient large-scale yield prediction. Deep learning (DL) (Goodfellow et al., 2016), a widely used ML method, has the advantage of leveraging a large amount of RS data by automatically learning features from big datasets. This removes the need for manual feature engineering, enabling the model to capture complex patterns and relationships between crop environmental factors and crop yield (He et al., 2016). Many studies applying DL to crop yield prediction have shown its advantages in performance improvement. You et al. (2017) introduced a scalable approach to crop yield prediction using modern representation learning rather than traditional hand-crafted features, achieving impressive results. Wang et al. (2018) demonstrated soybean yield predictions in Argentina and obtained reliable results in Brazil using transfer learning, even with limited data. Ma et al. (2021a) developed a county-level corn yield prediction model using a Bayesian Neural Network across 12 states in the Corn Belt, showing the potential for large-scale corn yield prediction tasks. However, despite the popularity of DL models in corn yield prediction tasks, they have certain limitations. First, DL models require a large amount of data to learn the relationships between inputs and outputs. Additionally, DL models often function as a black box, lacking reliable interpretability of the prediction results. Specifically, crop growth involves complex biological processes among soil, weather, and corn, which are challenging for DL models to fully capture. Therefore, how to leverage advantages of process-based models and DL models to address these challenges has become an important research topic.
@@
Knowledge-guided machine learning (KGML) (Karpatne et al., 2022) aims to integrate scientific knowledge in ML frameworks to achieve better performance, scientific consistency, and explainability of results. KGML involves three main approaches: Knowledge-Guided Learning, Knowledge-Guided Architecture, and Knowledge-Guided Pretraining. Knowledge-Guided Learning incorporates scientific knowledge, such as physical laws, into the learning algorithm by modifying the loss function. It guides the optimization process toward scientifically consistent and generalizable solutions (Daw et al., 2022; Bao et al., 2021). Knowledge-Guided Architecture embeds scientific knowledge directly into the structure of the ML model. By designing knowledge-guided neural network architectures, researchers can enable models to automatically capture physical properties or conservation laws, which improves both interpretability and robustness (Dugdale et al., 2017; Luo et al., 2023). Knowledge-Guided Pretraining uses simulated data or self-supervised learning tasks to pretrain ML models. By initializing model parameters using data that encodes scientific principles, this method reduces the need for large amounts of observational data and improves performance in data-scarce scenarios (Licheng et al., 2022; Chen et al., 2023). Building on these approaches, KGML has been increasingly used in various domains, such as agriculture, hydrology, and biology, enabling improved data-driven decision-making and predictive modeling.
@@
Recently, many studies have used KGML to combine process-based and ML models for crop yield predictions (Kimball et al., 2023; Burroughs et al., 2023). Since process-based models already provide insights into the crop growth process and generate high-quality simulated data, Knowledge-Guided Pretraining is the most popular approach (He et al., 2023; Yang et al., 2023). These studies aim to identify patterns between agricultural features and crop yields from large datasets using ML, while leveraging data simulated by process-based models to provide interpretability. Specifically, intermediate variables in crop growth generated by process-based models are used to train ML models, allowing the ML models to simulate specific parts of the crop growth process. For example, He et al. (2023) integrated physical knowledge from process-based models and enabled the model to adjust for temporal data shifts over the years. Yang et al. (2023) used both historical and in-season RS data to better predict key agricultural variables such as grain yield and carbon cycling processes. Yang et al. (2024) presented a knowledge-guided computer vision framework designed to monitor and simulate the growth of individual strawberry fruits, which can predict growth at the individual fruit level by integrating image-based data and dynamic simulations.
Although these KGML models have made significant progress in crop yield prediction, all existing KGML-based studies on crop yield predictions have overlooked the critical role of soil, particularly how soil moisture contributes to corn growth and affects yield. Soil moisture is influenced by temperature, precipitation, and other environmental factors and directly affects crop growth. As a key indicator of crop water supply, low soil moisture causes water stress, which limits corn growth and reduces yields (Unganai and Kogan, 1998). Unlike most RS data, such as NDVI, which only reflects the growth status of plants at a specific moment, soil moisture has a cumulative impact on yield over time. Nonetheless, a significant challenge for ML models is the limited availability of soil moisture data. For example, the widely used SPL4SMGP.007 SMAP soil moisture dataset (Reichle et al., 2022) only covers the period from 2015 onward, which limits the ability of ML models to learn from extensive historical data and impacts model performance. Therefore, we present a Knowledge-Guided Machine Learning with Soil Moisture (KGML-SM) framework that explicitly incorporated the soil moisture impact in the ML model and provided interpretability. The model consisted of two major components: a Weather-to-Soil (W2S) encoder modeling the relationship between weather inputs and soil moisture outputs, and an attention module (Vaswani, 2017) integrating features from multiple sources and dynamically weighting their influences on corn yield across time and space. To train this model, we adopted a two-step process. First, we used the APSIM model to generate a simulated field-level dataset for pretraining. Then, the model was finetuned using a county-level dataset from Google Earth Engine (GEE) (Gorelick et al., 2017) and United States Department of Agriculture (USDA) National Agricultural Statistics Service (NASS).
Furthermore, we conducted research into the relationship between soil moisture and corn yield prediction in drought-affected areas, where drought often reduced yields and caused overestimation (Cooper et al., 2014). Researchers have long been interested in studying the relationship between soil moisture, drought, and corn yield prediction (Ines et al., 2013; Bushong et al., 2016; Mladenova et al., 2017). However, these studies either had a narrow scope, or did not utilize this connection for model optimization. we designed a drought-aware loss function to further enhance model performance under drought conditions and mitigate overestimation problem. Finally, based on the relationship between soil moisture and corn yield prediction, we provided explanations for the prediction errors of the KGML-SM model and offered directions for future model optimization. Our research aimed to improve the accuracy and interpretability of corn yield predictions, ultimately contributing to more resilient agricultural systems under extreme weather variability.
Our main contributions are as follows:
  • 1.
    This paper introduced the KGML-SM framework. This framework integrated process-based and ML models for corn yield prediction while explicitly incorporating the influence of soil moisture.
  • 2.
    A drought-aware loss function was designed. This function enhanced model performance under drought conditions and mitigated overestimation.
  • 3.
    Based on the relationship between drought, soil moisture, and corn yield prediction, we provided interpretability for the prediction errors of the KGML-SM model and offered directions for future model optimization.
@@ 2 Data acquisition
In this study, we developed two datasets for corn yield prediction. The first dataset was a field-level dataset generated using APSIM (McCown et al., 1996) and used for pretraining. The second dataset was a county-level dataset derived from GEE (Gorelick et al., 2017) and USDA NASS and used for finetuning. The workflow was divided into three main steps: simulation, pretraining, and finetuning. A summary of all public datasets used in this study is provided in Table 1. In this section, we first introduce the study area (Sec. 2.1); then, we provide details of the APSIM field-level dataset (Sec. 2.2); finally, we describe the construction of the GEE county-level dataset (Sec. 2.3).
@@
MODIS MCD43A4 c Vegetation indices (Schaaf and Wang, 2015)
@@
SMAP c Soil moisture data (Reichle et al., 2022)
@@ 2.1 Study area
Our research focused on corn yield prediction across the U.S. Corn Belt. Twelve states are selected as our study area, including North Dakota, South Dakota, Minnesota, Wisconsin, Iowa, Illinois, Indiana, Ohio, Missouri, Kansas, Nebraska, and Michigan. These states are crucial agricultural states in the U.S., known for their significant contributions to corn production. We generated a five-year average yield map (Fig. 1) for these twelve states and considered them well suited for corn yield prediction research.
@@ 2.2 APSIM field-level dataset
In this section, we explain how APSIM was used to generate a field-level simulated dataset for model pretraining. We used data from the Iowa Environmental Mesonet (IEM) (Herzmann et al., 2004), a platform developed by Iowa State University that provides U.S. agricultural and environmental data. We extracted all available station data from the IEM, which included data from thousands of stations across 12 states from 1980 to 2023 (Table 2).
@@
State IL IN IA KS MI MN MO NE ND OH SD WI
Station number 120 89 113 150 140 140 130 128 79 100 114 147
2.2.1 Input and output of APSIM simulation
The APSIM model simulates corn yield based on four weather data inputs: maximum and minimum temperature (Tmax and Tmin), precipitation (PPT), and radiation (Radn). The simulation also requires the station’s latitude, longitude, and year. Some management parameters are also necessary, including start and end of sowing window, plant population, fertilizer amount, and initial soil water, etc (Sec. 3.3.1). Tmax and Tmin influence the rate of plant development, and affect processes like photosynthesis and respiration. PPT are essential for modeling soil moisture levels, which directly affect water availability for crops. Radn is a critical factor in photosynthesis, as it provides the energy needed for plant growth. These weather data of IEM stations are available from the IEM website (Herzmann et al., 2004). The APSIM model takes weather data as input to simulate root zone soil moisture (SM_rootzone), surface soil moisture (SM_surface), and corn yield. SM_rootzone is crucial for corn growth as it directly affects water availability for uptake, influencing plant development and yield. SM_surface plays a key role in seed germination and early growth stages.
2.2.2 Variables summary in field-level dataset for pretraining
We calculated a five-year historical average yield based on the simulated corn yield to establish a local baseline for model accuracy improvement. We also included the prediction year and location (longitude and latitude) of IEM stations in the field-level dataset. Because soil properties and topography vary across space and time, year and location can help the ML model capture this variation. Then we combined the weather data, simulated soil moisture and corn yield, along with other variables to construct the APSIM field-level dataset (Table 3).
Table 3: Summary of variables comprising the APSIM field-level dataset, including inputs and simulated outputs